OpenStudy (marinedude42):

What is the slope of the line given by the equation y = -5x? Enter your answer as an integer or fraction in lowest terms.

OpenStudy (anonymous):

y= mx + b; where m = the slope of the graph.

OpenStudy (texaschic101):

Martymar is correct....in y = mx + b form, the m is the slope. I know your equation says y = -5x, but look at it like this : y = -5x + 0....now what number is in the m position ?

OpenStudy (marinedude42):

In what direction is the line containing the points (-4, -6) and (-2, 8) going? A. up and to the right B. vertical C. horizontal D. down and to the right

OpenStudy (marinedude42):

d right?

OpenStudy (texaschic101):

first we find the slope by using the slope formula slope = (y2 - y1) / (x2 - x1) (-4,-6) x1 = -4 and y1 = -6 (-2,8) x2 = -2 and y2 = 8 time to sub slope = (8 - (-6) / (-2 - (-4) slope = (8 + 6) / (-2 + 4) slope = 14/2 slope = 7/1 slope is positive meaning it goes up and to the right

OpenStudy (anonymous):

A

OpenStudy (marinedude42):

can you guys help me with my apex please?